When 25. 0 mL of 0. 25 mol /L LiOH and 25. 0 mL of 0. 25 are mol/l HCl mixed together, the temperature warms to 15.8 °C What is the molar enthalpy of neutralization for LiOH?
1
Expert's answer
2020-11-20T08:09:36-0500

"LiOH + HCl \to LiCl + H_2O"


Number of moles of LiOH = 25/1000L × 0.25moles = 0.00625mol .


Total volume of solution = 25ml + 25ml = 50ml


mass of solution = volume × density of water(since the solution is a water solution) = 50ml × 1g/ml = 50g


Enthalpy of Neutralization + heat lost = 0


n∆H + mc∆T = 0


0.00625(∆H) + 50(4.184)(15.8) = 0

0.00625(∆H) = -3305.36

∆H = -528857.6 j/mol

∆H = -529 kJ/mol